OpenStudy (anonymous):

A new company will reease 3 product A,B,C.20% of buyer will buy A, 55% will buy B and 25% will buy C. People that buy A will buy 1.5 units,B 2.5 and C 0.75.What % of total sales will be product of B?

OpenStudy (anonymous):

can i get a detailed explanation for this?

OpenStudy (aravindg):

First assume the number of customersbe 100. 20 buys A, 55 buys B and 25 buys C. ie 55 out of 100 buys B. Now the amount bought by 55 people will be 2.5 times 55.

OpenStudy (anonymous):

the total shoppers are 20% + 55% + 25% (100%) = 0.2 + 0.55 + 0.25 = 1 the total purchases are 0.2*(1.5)+0.55*(2.5)+0.25*(0.75) = 0.3+1.375+0.1875 =1.8625 so total sales will not be 100%, but 186% relative to customers because A, B all purchase multiple units. (they outweigh the customers C which buy less than one.) out of the 186% total sales, B will buy 55% x 2.5 units = 137.5% of the total sales. when normalizing for 186% total sales, we divide by 1.86: 137.5% of 186% is equal to 73.92% of 100% . The product B accounts for 73.92% of all sales.

OpenStudy (aravindg):

Similarly calculate amount bought by A and C.

OpenStudy (anonymous):

thank u mathessentials and AravindG

OpenStudy (aravindg):

yw :) and Welcome to Openstudy!